# An NV- center in magnesium oxide as a spin qubit for hybrid quantum technologies

### v1

@@ Acknowledgements.
We thank Dr. F. Joseph Heremans, Prof. Shun Kanai, and Dr. Christian Vorwerk for many useful discussions. V.S acknowledges the support from the Maria Goeppert Mayer Named Fellowship, under the Laboratory Directed Research and Development (LDRD) funding from Argonne National Laboratory, provided by the Director, Office of Science, of the U.S. Department of Energy under Contract No. DE-AC02-06CH11357. J.D. acknowledges support from the Swedish e-science Research Centre (SeRC), the Knut and Alice Wallenberg Foundation through the WBSQD2 project (Grant No. 2018.0071), and the Swedish Research Council (VR) Grant No. 2022-00276. This work was supported by the Air Force Office of Scientific Research (AFOSR) through the CFIRE grant # FA95502310667. This work used several codes, including the WEST code whose development is supported by MICCoM, which is part of the Computational Materials Sciences Program funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences, and Engineering Division through Argonne National Laboratory. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award DDR-ERCAP0029604, resources of the University of Chicago Research Computing Center, and resources of the Argonne Leadership Computing Facility, a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC02-06CH11357.

### v2

@@ V Data availability
VI Competing interests
The Authors declare no Competing Financial or Non-Financial Interests.
VII Author contributions
V.S.: calculations and analyses; J.D.: ADAQ methodology and high-throughput screening; Y.J.: TDDFT and PL methodology; G.G.: designed and supervised the research. V.S. and G.G. wrote the manuscript. All authors contributed to discussions and finalizing the manuscript.
@@ Acknowledgements.
We thank Dr. F. Joseph Heremans, Prof. Shun Kanai, and Dr. Christian Vorwerk for many useful discussions. V.S acknowledges the support from the Maria Goeppert Mayer Named Fellowship, under the Laboratory Directed Research and Development (LDRD) funding from Argonne National Laboratory, provided by the Director, Office of Science, of the U.S. Department of Energy under Contract No. DE-AC02-06CH11357. J.D. acknowledges support from the Swedish e-science Research Centre (SeRC), the Knut and Alice Wallenberg Foundation through the WBSQD2 project (Grant No. 2018.0071), and the Swedish Research Council (VR) Grant No. 2022-00276. This work was supported by the Air Force Office of Scientific Research (AFOSR) through the CFIRE grant # FA95502310667. This work used several codes, including the WEST code whose development is supported by MICCoM, which is part of the Computational Materials Sciences Program funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences, and Engineering Division through Argonne National Laboratory. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award DDR-ERCAP0029604, resources of the University of Chicago Research Computing Center, and resources of the Argonne Leadership Computing Facility, a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC02-06CH11357. The ADAQ computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at the Swedish National Infrastructure for Computing (SNIC) at Tetralith, partially funded by the Swedish Research Council through grant agreement no. 2022-06725.